Three months ago, AXE stock was selling for $44.25 a share. At that time, you sold three put options on the stock with a strike price of $45 per share and an option price of $1.75 per share. The option expires today when the value of the stock is $42.50 per share. What is your net profit or loss on this investment? Ignore commissions and taxes. Round the answer to the nearest dollar. Enter as positive if a profit; enter as negative if a loss.
Answer #1

At expiration,

Value of Put Option = (45 - 42.50) = $2.50

So,

Return Rate = (2.50 - 1.75)/1.75

Net Profit = 3(2.50 - 1.75)

Net Profit = 3*1.25

Net Profit = $3.75